Find three consecutive whole numbers such that the product of the first and the second is 23 more than the third.
Answer by usyim88hk(158)

let x be the smallest number
(x)(x+1) = (x+2)+23
because the product of first two numbers is equal to the third number plus 23
so now solve it
x^2+x = x+25
x^2 = 25
x = 5
So the three numbers are 5,6 and 7